OpenStudy (anonymous):

Sergio has decided to diversify his investments in the following way: $4,000 in an account earning 3.2% simple interest $4,000 in a savings account earning 2.1% interest compounded annually $7,500 in a certificate of deposit earning 5% interest compounded quarterly How much total interest will Sergio earn on his investments at the end of 4 years?

OpenStudy (paki):

(4000*2.3/100)= 128 $... for 1 year... for 4 years... 128*4= 512$ (4000*2.1/100)= 84 $... for 1 year... for 4 years... 84*4= 336$ (7500*5/100)= 375 $... for 1 year... for 4 years... 375*4= 1500$ now sum up all the amounts... 512+336+1500= 2348$